I have a cookie jar with 10 cookies – 3 are chocolate chip, 2 are oatmeal, and 5 are sugar cookies. What is the probability that I get a chocolate chip cookie on my next try if I’ve already pulled out 2 oatmeal and 2 sugar cookies?
Answer #1

Since you already pulled out 2 oatmeal and 2 sugar cookies so number of cookies remaining is 10 -2-2 = 6. Out of 6 cookies, 3 are chocolate chip cookie so

P(chocolate chip cookie | pulled out 2 oatmeal and 2 sugar cookies) = 3 / 6 = 0.50

Answer: 0.50